# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v2

@@ 4.1 Gaia Challenge mocks
Figure 1: Upper left: 3D dark matter density profile for the PlumCoreOM mock galaxy, bands denote the 95% CL regions from the median for different numbers of stellar tracers including both LOS and PM velocities. The results from 2021MNRAS.501..978R for Agama $f(E,L)$ models, and from 2021MNRAS.505.5686C for GravSphere1.5 (including LOS VSPs) for 1,000 tracers with their respective median values with 95% CL errors are shown for comparison. The red line denotes the true analytical solution from which the mock was generated. The gray, dashed line corresponds to the projected half-light radius ($R_{1/2})$. Upper right: Same as the left hand side but for the PlumCuspOM model. Lower left: Symmetrized anisotropy profile for the PlumCoreOM model. The red line again denotes the true solution from which the mock was generated, corresponding to an Osipkov-Merritt profile. Lower right: Same as the left hand side but for the PlumCuspOM model.
@@
Figure 2: Same as Fig. 1 but only with LOS tracers. Upper left: GravSphere2 recovery for the PlumCoreOM mock 3D mass density. Upper right: Same as the left but for the PlumCuspOM model. Lower left: Symmetrized anisotropy profile for the PlumCoreOM model. Lower right: Same as the left but for the PlumCuspOM model.
@@
Figure 3: Same as Fig. 2 but for 10,000 tracers in the PlumCoreOM mock, including in the comparison with other models. We have included the DiscreteJAM results from 2021MNRAS.501..978R. Left: Mass density profiles. Blue band shows our results when Gaussianity is assumed and fourth-order moments are excluded. For comparison, the inset shows instead bands obtained from the posterior distribution of likelihoods when computing extrema of all models falling within the 68th percentile from the maximum likelihood (see text for details). Using this definition, there is a wide range of equally good-fitting models that encompasses the true solution, while the posterior becomes biased due to a relative abundance of cuspier models in the allowable hypervolume. Right: Same as left, but for symmetrized anisotropy profile.
@@
Figs. 1 and 2 show the results for the mass density and anisotropy recovery of the two models considered, with and without the inclusion of PMs (respectively). These are shown as the regions in the posterior spanning 95% of the distribution centered at the median values of the profiles. For reference, we have included the results of GravSphere1.5 (originally from 2021MNRAS.505.5686C), [^18] Note that 2021MNRAS.505.5686C only present their results for LOS tracers. We have, however, also included their (until now unpublished) results including also PMs. the most recent version of GravSphere  until now, and the results presented in 2021MNRAS.501..978R for the Agama $f(E,L)$ distribution function-based models, which was in most cases the best-performing method (or amongst the best-performing ones) of the multiple ones considered in 2021MNRAS.501..978R. This distribution function method is part of the widely-used Agama library (2019MNRAS.482.1525V). These cases shown for comparison use 1,000 tracers.
@@
While GravSphere1.5 shows better performance than GravSphere1 in most cases for this mock sample (cf. 2021MNRAS.501..978R), it still suffers from significant bias, particularly for the cored model, where densities are systematically underestimated beyond $\sim R_{1/2}$, even when PMs are included (Figs. 1 and 2). GravSphere2, on the other hand, is able to recover the true solutions within the 95% confidence-level (CL) region from the median within $0.1\lesssim r/R_{1/2}\lesssim 10$ that encloses the vast majority of the tracers. The only exception to this concerns cases with limited data, where we can see that for the core model with 100 LOS tracers (Fig. 2), the model underestimates the anisotropy beyond $\sim 2R_{1/2}$, but is nonetheless able to recover the density everywhere at the specified confidence. For the 1,000-tracer case, the true solution also lies at the edge of the $95\%$ CL region beyond $R_{1/2},$ though these mild biases disappear when the number of tracers is increased or PMs are included, which is not the case with GravSphere1.5. The performance is otherwise excellent and in all cases competitive with sophisticated modeling approaches like Agama, which shows a similar performance in most cases. We even see some instances where GravSphere2 shows less bias than Agama (e.g. Fig. 1, upper left) at larger radii, and reduced errors overall. [^19] We emphasize, however, that we are referring to the Agama model applied in 2021MNRAS.501..978R. This means that the biases that we see strictly speaking apply only to the phase-space distribution function parametrization adopted by these authors.
@@
Another illustrative case is shown in Fig. 3, where we compare the core model with 10,000 LOS tracers. Here all the models considered by 2021MNRAS.501..978R were biased in either the density or anisotropy recovery, with GravSphere1.5, GravSphere1, and Agama showing a degree of bias in both. GravSphere2, on the other hand, is able to obtain a 95% CL recovery everywhere throughout. Models that do not capture deviations of Gaussianity in this case suffer particularly strong bias. To illustrate this, we included the results of DiscreteJAM axisymmetric Jeans models, which assume Gaussian velocity PDFs. We then compared this with the result of imposing a Gaussian PDF in our models and not considering higher moments. This yields a similar result in both cases with a clearly biased model showing significantly greater slopes and uncertainty. GravSphere1.5 and GravSphere1, with VSPs, and Agama, which also allows more generality in the velocity PDF modeling, are able to achieve better recovery than the other models considered in 2021MNRAS.501..978R, but still suffer from a degree of bias across much of the radial range shown ($0.25\lesssim r/R_{1/2}\lesssim 4$). We also note that the anisotropy priors imposed by 2021MNRAS.505.5686C on GravSphere1.5, which we also impose, mean that our models are barred from large negative anisotropies and are close to isotropic at the center. Similarly, Agama models are fixed to be purely radially anisotropic at large radii. This partly explains why these models tend to be more tightly constrained towards the edges of our anisotropy plots, but, at least for GravSphere2, our results are otherwise insensitive to these priors.
@@
Our results show that the biases we observe can be overcome in GravSphere2 by including a full treatment of higher moments via the extended Jeans equation (rather than using VSPs, which can be prone to bias themselves when determining them – see the discussion at the end of Sec. 2.1). The inclusion of these higher moments also removes negative-moment models at fourth order, while we extract more information from the data than GravSphere1.5 by fitting the individual velocities with our generalized PDF modeling. For distribution function-based models, the challenge consists in creating sufficiently generalized functional forms so as to eliminate any biases. On the other hand, it is apparent that models which assume Gaussian velocity PDFs will be prone to biased recovery if the velocity distributions of the system, as in the case of Fig. 3, are intrinsically non-Gaussian.
@@
In Fig. 4, we show the recovery of the logarithmic density slope for the Gaia Challenge mocks. In all cases, GravSphere2 is able to recover these within its 95% confidence intervals within $2R_{1/2},$ and in most cases everywhere in the plot and beyond ($0.1\lesssim r/R_{1/2}\lesssim 10$). For the 1,000-tracer case, only GravSphere2 recovers the true solution everywhere within the plots in all mocks. GravSphere1.5 systematically overestimates the steepness of the slope below $R_{1/2}$ for the LOS core model. Agama, on the other hand, is able to recover the slope profile well for all cases except for the core model with PM data beyond $\gtrsim R_{1/2}$, where moderate bias is observed (and also with higher uncertainties towards the center at $r\lesssim R_{1/2}$). The only case where a mild degree of bias is observed for GravSphere2 is for the LOS-only cusp model beyond $2R_{1/2}$, where the true slope profile narrowly falls outside the 95% CL region. In this case, we verified that this was due to the coreNFWtides model slightly overestimating slopes in these regions. If we adopt instead an $\alpha\beta\gamma$ model (Eq. 33), which was used to generate the mocks (and also adopted for the Agama results), the profile was recovered at all radii without any bias. This test is shown in Appendix A. [^20] Note that it is not very surprising that biases will be naturally reduced when assuming the same mass model as the one used to generate the mock data. For real data, of course, we cannot know what the true underlying model is, so it is prudent to explore a couple of different options where possible.
@@
To better understand the reason for the biases observed with other methods, we also show in Fig. 3 (inset) the results of taking the maxima and minima of models falling within the 68th percentile from the maximum likelihood taken from the posterior distribution. This is done as follows: (i) we compute the posterior of the likelihoods; (ii) determine the limiting value corresponding to the 32nd percentile, so that 68% of models are equal or above this likelihood (the 100th percentile corresponding to the maximum likelihood); (iii) for each radius, we calculate the minima and maxima of all models with likelihoods greater or equal than our threshold value. This approach differs from the classical Bayesian posterior and we have chosen it as a measure to define the space of models which can be said to be ‘equally good-fitting’. It is more reminiscent of frequentist statistics and we intend to explore it in more detail in upcoming works as a means of testing prior and posterior hypervolume shape dependence in models.
@@
The latter result leads us to propose a diagnostic that may be important to better understand differences between quoted results for e.g. the slopes and density profiles of dwarf galaxies, an active topic of debate and discussion in the literature (e.g. 2018MNRAS.481..860R; 2020ApJ...904...45H; 2024ApJ...970....1V; 2025A&A...699A.347A). If bounds from the posterior confidence intervals show significantly less spread (i.e. appear more aggressive) than the ones from taking the extremal values of models of comparable likelihoods (with our reference test being the 68th percentile from the maximum), then this indicates that the results are dominated by the shape of the posterior hypervolume, which is sensitive to priors and model parameterizations. Consequently, one should on these occasions proceed with caution when interpreting the results, which may – without careful consideration – overstate the capacity of the analysis to truly distinguish among competing hypotheses. Conversely, if one does not find such a discrepancy, then it is indicative of the fact that the favored models are indeed better-fitting than the disfavored ones and not merely more abundant in the posterior hypervolume.
As noted in Sec. 3.3, however, the priors we adopt in our models are generally very broad compared to the results favored by the data, which we show explicitly in Appendix B. This shows that the posterior hypervolume dependence that we observe with our diagnostic is not limited to cases where the posterior parameter space spans the entire prior volume, but more generally to degenerate regions allowed by the data within sub-volumes of the prior space, especially when the true solution lies at the edge of the allowable hypervolume. This can be conditioned by a biased likelihood, which is partly at play in our example assuming Gaussianity, and also by the prior itself (the hypervolume region is the intersection of the models with allowable likelihoods that lie in the prior space). Also, even in cases where the model correctly recovers the true solution, meaning that the hypervolume is not skewed, a large spread when comparing with our diagnostic can indicate that the confidence of the model is overstated by posterior statistics (i.e. that the posterior hypervolume happens to have more models clustering around the true solution). In this way, our diagnostic based purely on the likelihood does not attempt to compute posterior statistics based on the relative abundance of models, as with typical Bayesian approaches, but merely computes the spread on a given quantity allowable within a likelihood threshold, and as a result will not be conditioned by these potential biases.
@@
As noted in Sec. 3.3, we have also included marginalized plots of parameters in a public repository for all mocks considered in this study. We note that individual parameters are well constrained in most cases with respect to the prior limits, and even in cases where this is not the case, physical quantities of interest, like the mass density in regions of interest, are insensitive to these degeneracies (c.f. Appendix B).
Lastly, we have included binned dispersion and kurtosis profiles for reference in Appendix D. The binned profiles are generally consistent with those derived from our GravSphere2 mass models and show meaningful deviations from Gaussianity. This, once again, stresses the importance of using models which are able to capture these deviations.
@@ 4.2 Simulated dwarf galaxies
Figure 5: Upper left: GravSphere2 recovery of the 3D dark matter density profile for the Fornax-like simulated galaxy, where the bands denote the 95% CL regions for different numbers of stellar tracers, as marked, including only LOS velocities. The total sample of bound tracers is 34,158 (gray band). The purple points correspond to the median values with 95% CL errors from 2025arXiv250418617T using all LOS tracers with GravSphere1. The red, solid line corresponds to the true density profile from the simulation (smoothened with a Savitzky-Golay filter). The gray, dashed line corresponds to the projected half-light radius ($R_{1/2}$). Upper right: Symmetrized anisotropy profile results. The red points denote the binned 3D anisotropies obtained from bins of equal numbers of tracers and directly fitting the generalized velocity PDF from Eqs. (25) and (2.2) along the three spherical-coordinate directions. Lower left: Same as Upper Left but including both LOS and PM velocities. Lower right: Same as Upper Right but with both LOS and PM velocities.
@@
Figure 7: Left: GravSphere2 recovery of the mass profile for the Fornax simulated galaxy with only LOS velocities for low stellar tracer numbers (10, 25, 50). The bands display the 95% CL values. We show the results obtained using the mass estimators from 2009ApJ...704.1274W; 2010MNRAS.406.1220W; 2018MNRAS.481.5073E. Right: Same, but for the PlumCoreOM mock with 100 tracers and Gaussian PDF binning.
Fig. 5 shows the GravSphere2 recovery of the mass density and anisotropy profiles for the Fornax-like simulated dwarf galaxy with and without PMs in addition to the LOS velocities, whilst Fig. 6 shows the recovery of the logarithmic density slope profiles for these models. The recovery of the density, anisotropy and slope using all available bound tracers with LOS velocities is excellent. Using only a small subset of these still yields very good recovery that is further improved by including PMs, leading to a comparable result to the Gaia Challenge equilibrium mocks from Sec. 4.1, and with true values falling everywhere ($0.1<r/R_{1/2}<10$) within the 95% CL regions. This, on the whole, suggests that these simulated dwarf galaxies do not show significant departures from dynamical equilibrium, spherical symmetry, or the assumption of no rotation. Comparing our results for the all-tracer LOS model with GravSphere1 from 2025arXiv250418617T (purple points in Fig. 5, upper left), GravSphere2 is significantly less biased than GravSphere1 for the density recovery, whilst also showing substantially smaller uncertainties overall (particularly towards the central region). Whilst 2025arXiv250418617T show that their biases can be substantially reduced by removing VSPs from their fits, this is done at the expense of higher uncertainties. They also obtained less biased results when weakening the monotonicity prior constraint on their non-parametric mass model, though some bias still remained (see their Fig. D1, also Fig. 8, where their biases are most apparent).
Lastly, we have included the projected binned velocity dispersion and kurtosis profiles in Appendix E. These are consistent with the ones inferred from our models using individual velocities and are close to isotropic, not showing large deviations from Gaussianity. The isotropy of the distribution is most evident from the right hand side of Fig. 5. This is less evident in 2025arXiv250418617T, but our binning procedure effectively removes the stochastic noise by averaging over large numbers of tracers, revealing a profile which is close to isotropic.

### v3

@@ 4.1 Gaia Challenge mocks
Figure 1: Top-left: 3D dark matter density profile for the PlumCoreOM mock galaxy, bands denote the 95% CL regions from the median for different numbers of stellar tracers including both LOS and PM velocities. The results from Read et al. (2021) for Agama $f(E,L)$ models, and from Collins et al. (2021) for GravSphere1.5 (including LOS VSPs) for 1,000 tracers with their respective median values with 95% CL errors are shown for comparison. The red line denotes the true analytical solution from which the mock was generated. The gray, dashed line corresponds to the projected half-light radius ($R_{1/2})$. Top-right: Same as the left hand side but for the PlumCuspOM model. Bottom-left: Symmetrized anisotropy profile for the PlumCoreOM model. The red line again denotes the true solution from which the mock was generated, corresponding to an Osipkov-Merritt profile. Lower right: Same as the left hand side but for the PlumCuspOM model.
@@
Figure 2: Same as Fig. 1 but only with LOS tracers. Top-left: GravSphere2 recovery for the PlumCoreOM mock 3D mass density. Top-right: Same as the left but for the PlumCuspOM model. Bottom-left: Symmetrized anisotropy profile for the PlumCoreOM model. Bottom-right: Same as the left but for the PlumCuspOM model.
@@
Figure 3: Same as Fig. 2 but for 10,000 tracers in the PlumCoreOM mock, including in the comparison with other models. We have included the DiscreteJAM results from Read et al. (2021). Left: Mass density profiles. Blue band shows our results when Gaussianity is assumed and fourth-order moments are excluded. For comparison, the inset shows instead bands obtained from the posterior distribution of likelihoods when computing extrema of all models falling within the 68th percentile from the maximum likelihood (see text for details). Using this definition, there is a wide range of equally good-fitting models that encompasses the true solution, while the posterior becomes biased due to a relative abundance of cuspier models in the allowable hypervolume. Right: Same as left, but for symmetrized anisotropy profile.
@@
Figs. 1 and 2 show the results for the mass density and anisotropy recovery of the two models considered, with and without the inclusion of PMs. These are shown as the regions in the posterior spanning 95% of the distribution centered at the median values of the profiles. For reference, we included the results of GravSphere1.5 (originally from Collins et al. 2021), [^20] Note that Collins et al. (2021) only present their results for LOS tracers. We have, however, also included their (until now unpublished) results including also PMs. the most recent version of GravSphere  until now, and the results presented in Read et al. (2021) for the Agama $f(E,L)$ distribution function-based models, which was in most cases the best-performing method (or amongst the best-performing ones) of the multiple ones considered in Read et al. (2021). This distribution function method is part of the widely-used Agama library (Vasiliev, 2019). These cases shown for comparison use 1,000 tracers.
@@
While GravSphere1.5 shows better performance than GravSphere1 in most cases for this mock sample (cf. Read et al. 2021), it still suffers from significant bias, particularly for the cored model, where densities are systematically underestimated beyond $\sim R_{1/2}$, even when PMs are included (Figs. 1 and 2). GravSphere2, on the other hand, is able to recover the true solutions within the 95% confidence-level (CL) region from the median within $0.1\lesssim r/R_{1/2}\lesssim 10$ that encloses the vast majority of the tracers. The only exception to this concerns cases with limited data, where we can see that for the core model with 100 LOS tracers (Fig. 2), the model underestimates the anisotropy beyond $\sim 2R_{1/2}$, but is nonetheless able to recover the density everywhere at the specified confidence. For the 1,000-tracer case, the true solution also lies at the edge of the $95\%$ CL region beyond $R_{1/2},$ though these mild biases disappear when the number of tracers is increased or PMs are included, which is not the case with GravSphere1.5. The performance is otherwise excellent and in all cases competitive with sophisticated modeling approaches such as Agama, which shows a similar performance in most cases. We even see some instances where GravSphere2 shows less bias than Agama (e.g., Fig. 1, upper left) at larger radii, and reduced errors overall. [^21] We emphasize, however, that we are referring to the Agama model applied in Read et al. (2021). This means that the biases that we see strictly speaking apply only to the phase-space distribution function parametrization adopted by these authors.
@@
Another illustrative case is shown in Fig. 3, where we compare the core model with 10,000 LOS tracers. Here all the models considered by Read et al. (2021) were biased in either the density or anisotropy recovery, with GravSphere1.5, GravSphere1, and Agama showing a degree of bias in both. GravSphere2, on the other hand, is able to obtain a 95% CL recovery everywhere throughout. Models that do not capture deviations of Gaussianity in this case suffer particularly strong bias. To illustrate this, we included the results of DiscreteJAM axisymmetric Jeans models, which assume Gaussian velocity PDFs. We then compared this with the result of imposing a Gaussian PDF in our models and not considering higher moments. This yields a similar result in both cases with a clearly biased model showing significantly greater slopes and uncertainty. GravSphere1.5 and GravSphere1, with VSPs, and Agama, which also allows more generality in the velocity PDF modeling, are able to achieve better recovery than the other models considered in Read et al. (2021), but still suffer from a degree of bias across much of the radial range shown ($0.25\lesssim r/R_{1/2}\lesssim 4$). We also note that the anisotropy priors imposed by Collins et al. (2021) on GravSphere1.5, which we also impose, mean that our models are barred from large negative anisotropies and are close to isotropic at the center. Similarly, Agama models are fixed to be purely radially anisotropic at large radii. This partly explains why these models tend to be more tightly constrained toward the edges of our anisotropy plots, but, at least for GravSphere2, our results are otherwise insensitive to these priors.
@@
Our results show that the biases we observe can be overcome in GravSphere2 by including a full treatment of higher moments via the extended Jeans equation (rather than using VSPs, which can be prone to bias themselves when determining them; see the discussion at the end of Sect. 2.1). The inclusion of these higher moments also removes negative-moment models at fourth order, while we extract more information from the data than GravSphere1.5 by fitting the individual velocities with our generalized PDF modeling. For distribution function-based models, the challenge consists in creating sufficiently generalized functional forms so as to eliminate any biases. On the other hand, it is apparent that models which assume Gaussian velocity PDFs will be prone to biased recovery if the velocity distributions of the system, as in the case of Fig. 3, are intrinsically non-Gaussian.
@@
In Fig. 4, we show the recovery of the logarithmic density slope for the Gaia Challenge mocks. In all cases, GravSphere2 is able to recover these within its 95% confidence intervals within $2R_{1/2},$ and in most cases everywhere in the plot and beyond ($0.1\lesssim r/R_{1/2}\lesssim 10$). For the 1,000-tracer case, only GravSphere2 recovers the true solution everywhere within the plots in all mocks. GravSphere1.5 systematically overestimates the steepness of the slope below $R_{1/2}$ for the LOS core model. Agama, on the other hand, is able to recover the slope profile well for all cases except for the core model with PM data beyond $\gtrsim R_{1/2}$, where moderate bias is observed (and also with higher uncertainties toward the center at $r\lesssim R_{1/2}$). The only case where a mild degree of bias is observed for GravSphere2 is for the LOS-only cusp model beyond $2R_{1/2}$, where the true slope profile narrowly falls outside the 95% CL region. In this case, we verified that this was due to the coreNFWtides model slightly overestimating slopes in these regions. If we adopt instead an $\alpha\beta\gamma$ model (Eq. 33), which was used to generate the mocks (and also adopted for the Agama results), the profile was recovered at all radii without any bias. This test is shown in Appendix A. [^22] Note: it is not very surprising that biases will be naturally reduced when assuming the same mass model as the one used to generate the mock data. For real data, of course, we cannot know what the true underlying model is, so it is prudent to explore a couple of different options where possible.
@@
To better understand the reason for the biases observed with other methods, we also show in Fig. 3 (inset) the results of taking the maxima and minima of models falling within the 68th percentile from the maximum likelihood taken from the posterior distribution. These results were achieved as follows: (i) we compute the posterior of the likelihoods; (ii) determine the limiting value corresponding to the 32nd percentile, so that 68% of models are equal or above this likelihood (the 100th percentile corresponding to the maximum likelihood); (iii) for each radius, we calculate the minima and maxima of all models with likelihoods greater or equal than our threshold value. This approach differs from the classical Bayesian posterior and we have chosen it as a measure to define the space of models which can be said to demonstrate an ”equally good fit.” It is more reminiscent of frequentist statistics and we intend to explore it in more detail in upcoming works as a means of testing prior and posterior hypervolume shape dependence in models.
@@
The latter result leads us to propose a diagnostic that may be important to better understand differences between quoted results for, for instance, the slopes and density profiles of dwarf galaxies, an active topic of debate and discussion in the literature (e.g., Read et al. 2018; Hayashi et al. 2020; Vitral et al. 2024; Arroyo-Polonio et al. 2025). If bounds from the posterior confidence intervals show significantly less spread (i.e., appear more aggressive) than the ones resulting from the extremal values of models of comparable likelihoods (with our reference test being the 68th percentile from the maximum), this indicates that the results are dominated by the shape of the posterior hypervolume, which is sensitive to priors and model parameterizations. Consequently, one should on these occasions proceed with caution when interpreting the results, which might (without careful consideration) overstate the capacity of the analysis to truly distinguish among competing hypotheses. Conversely, if one does not find such a discrepancy, then it is indicative of the fact that the favored models are indeed better-fitting than the disfavored ones and not merely more abundant in the posterior hypervolume.
As noted in Sect. 3.3, however, the priors we adopt in our models are generally very broad compared to the results favored by the data, which we show explicitly in Appendix B. This shows that the posterior hypervolume dependence that we observe with our diagnostic is not limited to cases where the posterior parameter space spans the entire prior volume, but more generally to degenerate regions allowed by the data within sub-volumes of the prior space, especially when the true solution lies at the edge of the allowable hypervolume. This can be conditioned by a biased likelihood, which is partly at play in our example assuming Gaussianity, and also by the prior itself (noting the hypervolume region is the intersection of the models with allowable likelihoods that lie in the prior space). Also, even in cases where the model correctly recovers the true solution, meaning that the hypervolume is not skewed, a large spread when comparing with our diagnostic can indicate that the confidence of the model is overstated by posterior statistics (i.e., that the posterior hypervolume happens to have more models clustering around the true solution). In this way, our diagnostic based purely on the likelihood does not attempt to compute posterior statistics based on the relative abundance of models, as with typical Bayesian approaches, but merely computes the spread on a given quantity allowable within a likelihood threshold, and as a result will not be conditioned by these potential biases.
@@
As noted in Sect. 3.3, we also included marginalized plots of parameters in a public repository for all mocks considered in this study. We note that individual parameters are well constrained in most cases with respect to the prior limits, and even in cases where this is not the case, physical quantities of interest, such as the mass density in regions of interest, are insensitive to these degeneracies (c.f. Appendix B).
Lastly, we included binned dispersion and kurtosis profiles for reference in Appendix D. The binned profiles are generally consistent with those derived from our GravSphere2 mass models and show meaningful deviations from Gaussianity. This, once again, stresses the importance of using models which are able to capture these deviations.
@@ 4.2 Simulated dwarf galaxies
Figure 5: Top-left: GravSphere2 recovery of the 3D dark matter density profile for the Fornax-like simulated galaxy, where the bands denote the 95% CL regions for different numbers of stellar tracers, as marked, including only LOS velocities. The total sample of bound tracers is 34,158 (gray band). The purple points correspond to the median values with 95% CL errors from Tchiorniy and Genina (2025) using all LOS tracers with GravSphere1. The red, solid line corresponds to the true density profile from the simulation (smoothened with a Savitzky-Golay filter). The gray, dashed line corresponds to the projected half-light radius ($R_{1/2}$). Top-right: Symmetrized anisotropy profile results. The red points denote the binned 3D anisotropies obtained from bins of equal numbers of tracers and directly fitting the generalized velocity PDF from Eqs. (25) and (2.2) along the three spherical-coordinate directions. Lower left: Same as Upper Left but including both LOS and PM velocities. Lower right: Same as Upper Right but with both LOS and PM velocities.
@@
Figure 7: Left: GravSphere2 recovery of the mass profile for the Fornax simulated galaxy with only LOS velocities for low stellar tracer numbers (10, 25, 50). The bands display the 95% CL values. We show the results obtained using the mass estimators from Walker et al. (2009); Wolf et al. (2010); Errani et al. (2018). Right: Same, but for the PlumCoreOM mock with 100 tracers and Gaussian PDF binning.
Fig. 5 shows the GravSphere2 recovery of the mass density and anisotropy profiles for the Fornax-like simulated dwarf galaxy with and without PMs in addition to the LOS velocities, whilst Fig. 6 shows the recovery of the logarithmic density slope profiles for these models. The recovery of the density, anisotropy and slope using all available bound tracers with LOS velocities is excellent. Using only a small subset of these still yields very good recovery that is further improved by including PMs, leading to a comparable result to the Gaia Challenge equilibrium mocks from Sect. 4.1, and with true values falling everywhere ($0.1<r/R_{1/2}<10$) within the 95% CL regions. This, on the whole, suggests that these simulated dwarf galaxies do not show significant departures from dynamical equilibrium, spherical symmetry, or the assumption of no rotation. Comparing our results for the all-tracer LOS model with GravSphere1 from Tchiorniy and Genina (2025) (purple points in Fig. 5, upper left), GravSphere2 is significantly less biased than GravSphere1 for the density recovery, whilst also showing substantially smaller uncertainties overall (particularly toward the central region). Whilst Tchiorniy and Genina (2025) show that their biases can be substantially reduced by removing VSPs from their fits, this is done at the expense of higher uncertainties. They also obtained less biased results when weakening the monotonicity prior constraint on their nonparametric mass model, though some bias still remained (see their Fig. D1, also Fig. 8, where their biases are most apparent).
Lastly, we included the projected binned velocity dispersion and kurtosis profiles in Appendix E. These are consistent with the ones inferred from our models using individual velocities and are close to isotropic, not showing large deviations from Gaussianity. The isotropy of the distribution is most evident from the right hand side of Fig. 5. This is less evident in Tchiorniy and Genina (2025), but our binning procedure effectively removes the stochastic noise by averaging over large numbers of tracers, revealing a profile which is close to isotropic.